LINE SHAPES IN THE Q BRANCH OF THE ν_{4} BAND OF METHANE

\underline{D. CHRIS BENNER}, V MALATHY DEVI, Department of Physics, The College of William and Mary, Williamsburg, VA 23187-8795; MARY ANN H. SMITH, Science Directorate, NASA Langley Research Center, MS 401A, Hampton, VA 23681-2199; LINDA R. BROWN, Jet Propulsion Laboratory, California Institute of Technology, 4800 Oak Grove Dr., Pasadena, CA 91109-8099.

The Tropospheric Emission Spectrometer on the Earth Observing System Aura satellite measures methane in the Earth’s atmosphere. Some of the methane channels are located in the Q branch of the ν_{4} band of methane at 1300 cm^{−1}. Using laboratory spectra obtained with the National Solar Observatory’s Fourier transform spectrometer at the McMath-Pierce solar telescope, this Q branch was examined for line mixing and speed dependent line shapes for both self-broadening and air-broadening. The analysis also determined line widths, pressure-induced shifts and line mixing as functions of temperature. Comparisons are made to demonstrate the importance of considering line shape effects in satellite measurements of the Earth’s atmosphere.^{a}

aThe research described in this paper was carried out by the College of William and Mary, Langley Research Center and Jet Propulsion Laboratory, California Institute of Technology under contract with the National Aeronautics and Space Administration.